# Physics Informed Neural Network Code for 2D Transient Problems (PINN-2DT) Compatible with Google Colab

### v1

@@ 2 Introduction
The goal of this paper is to replace the functionality of the time-dependent solver we published using isogeometric analysis and fast alternating directions solver [5, 6, 7] with the Physics Informed Neural Network (PINN) python library that can be easily executed on Colab. The PINN proposed in 2019 by Prof. Karniadakis revolutionized the way in which neural networks find solutions to initial-value problems described using partial differential equations [1] This method treats the neural network as a function approximating the solution of the given partial differential equation $u(x)=PINN(x)$. After computing the necessary differential operators, the neural network and its appropriate differential operators are inserted into the partial differential equation. The residuum of the partial differential equation and the boundary-initial conditions are assumed as the loss function. The learning process involves sampling the loss function at different points by calculating the PDE residuum and the initial boundary conditions. The PINN methodology has had exponential growth in the number of papers and citations since its creation in 2019. It has multiple applications, from solid mechanics [15], geology [4], medical applications [11], and even the phase-field modeling of fracture [14]. Why use PINN solvers instead of classical or higher order finite element methods (e.g., isogeometric analysis) solvers? PINN/VPINN solvers have affordable computational costs. They can be easily implemented using pre-existing libraries and environments (like Pytorch and Google Colab). They are easily parallelizable, especially on GPU. They have great approximation capabilities, and they enable finding solutions to a family of problems. With the introduction of modern stochastic optimizers such as ADAM [3], they easily find high-quality minimizers of the loss functions employed.
@@ 5.1 Heat transfer
               dfdx(pinn, x, y, t) ** 2 - \
               dfdy(pinn, x, y, t) ** 2
@@ 5.2 Wave equations
$\frac{\partial u}{\partial n}=0\textrm{ }(x,y)\in\partial\Omega$ (18)
@@ 5.3 Thermal inversion
In this example, we aim to model the thermal inversion effect. The numerical results presented in this section are the PINN version of the thermal inversion simulation performed using isogeometric finite element method code [5] described in [9]. The scalar field $u$ in our simulation represents the water vapor forming a cloud. The source represents the evaporation of the cloud evaporation of water particles near the ground. The thermal inversion effect is obtained by introducing the advection field as the gradient of the temperature. Following [10] we define $\frac{\partial T}{\partial y}=-2$ for lower half of the domain ($y<0.5$), and $\frac{\partial T}{\partial y}=2$ for upper half of the domain ($y>0.5$).
@@
$\frac{\partial u(x,y,t)}{\partial t}+\left(b(x,y,t)\cdot\nabla\right)u(x,y,t)-\nabla\cdot\left(K\nabla u(x,y,t)\right)=f(x,y,t)\textrm{ }(x,y,t)\in\Omega\times(0,T]$ (24)
$\nabla u\cdot n=0\textrm{ in }\partial\Omega\times(0,T]$ (25)
$u(x,y,0)=u_{0}(x,y)\textrm{ in }\Omega\times 0$ (26)
@@
$\frac{\partial u(x,y,t)}{\partial t}+\frac{\partial T(y)}{\partial y}\frac{\partial u(x,y,t)}{\partial y}-0.1\frac{\partial u(x,y,t)}{\partial x^{2}}-0.01\frac{\partial u(x,y,t)}{\partial y^{2}}=f(x,y,t)\textrm{ }(x,y,t)\in\Omega\times(0,T]$ (28)
$\nabla u\cdot n=0\textrm{ in }\partial\Omega\times(0,T]$ (29)
$u(x,y,0)=u_{0}(x,y)\textrm{ in }\Omega\times 0$ (30)
@@ 5.4 Tumor growth
$\nabla u\cdot n=0\textrm{ in }\partial\Omega\times(0,T]$ (35)
$u(x,y,0)=u_{0}(x,y)\textrm{ in }\Omega\times 0$ (36)
@@
$-\frac{\partial D(x,y)}{\partial y}\frac{\partial u(x,y,t)}{\partial y}-D(x,y)\frac{\partial^{2}u(x,y,t)}{\partial x^{y}}-\rho u(x,y,t)\left(1-u(x,y,t)\right)=0$ (38)
[…]
@@
$\left(\frac{\partial u(x,y,t)}{\partial t}-\frac{\partial D(x,y)}{\partial x}\frac{\partial u(x,y,t)}{\partial x}-D(x,y)\frac{\partial^{2}u(x,y,t)}{\partial x^{2}}\right.$
$\left.-\frac{\partial D(x,y)}{\partial y}\frac{\partial u(x,y,t)}{\partial y}-D(x,y)\frac{\partial^{2}u(x,y,t)}{\partial x^{y}}-\rho u(x,y,t)\left(1-u(x,y,t)\right)\right)^{2}$ (41)

### v2

@@ 2 Introduction
The goal of this paper is to replace the functionality of the time-dependent solver we published using isogeometric analysis and fast alternating directions solver [8, 9, 10] with the Physics Informed Neural Network (PINN) python library that can be easily executed on Colab. The PINN proposed in 2019 by Prof. Karniadakis revolutionized the way in which neural networks find solutions to initial-value problems described using partial differential equations [1] This method treats the neural network as a function approximating the solution of the given partial differential equation $u(x)=PINN(x)$. After computing the necessary differential operators, the neural network and its appropriate differential operators are inserted into the partial differential equation. The residuum of the partial differential equation and the boundary-initial conditions are assumed as the loss function. The learning process involves sampling the loss function at different points by calculating the PDE residuum, the residuum of the initial conditions, as well as the residuum of the boundary conditions. The PINN methodology has had exponential growth in the number of papers and citations since its creation in 2019. It has multiple applications, from solid mechanics [18], geology [7], medical applications [14], and even the phase-field modeling of fracture [17]. PINN/VPINN solvers have affordable computational costs. They can be easily implemented using pre-existing libraries and environments (like Pytorch and Google Colab). They are easily parallelizable, especially on GPU. They have great approximation capabilities, and they enable finding solutions to a family of problems. There are several modern stochastic optimizers [3, 6, 4, 5], where the ADAM algorithm seems to be the most popular [3]. They easily find high-quality minimizers of the loss functions employed.
@@ 3 Physics Informed Neural Network for transient problems on the example of heat transfer problem
$u(x,y,t)=PINN(x,y,t)=A_{n}\sigma\left(A_{n-1}\sigma(...\sigma(A_{1}\begin{bmatrix}x\\ y\\ t\end{bmatrix}+B_{1})...+B_{n-1}\right)+B_{n}$ (7)
@@
$LOSS_{BC}(x,y,t)=\left(\frac{\partial PINN(x,y,t)}{\partial n}-0\right)^{2}$ (10)
@@ 5.1 Heat transfer
               dfdx(pinn, x, y, t, order=2)  - \
               dfdy(pinn, x, y, t, order=2)
@@ 5.2 Wave equations
$\frac{\partial u(x,y,t)}{\partial n}=0,\textrm{ }(x,y,t)\in\partial\Omega\times[0,T]$ (21)
Here $z(x,y)$ stands for the seabed and seashore topography, $g=9.81$ is the acceleration due to the Earth gravity, and $u_{0}(x,y)$ is the initial shape of the wave. We expand the wave equation by computing the partial derivatives.
@@
$u(x,y,t)=PINN(x,y,t)=A_{n}\sigma\left(A_{n-1}\sigma(...\sigma(A_{1}\begin{bmatrix}x\\ y\\ t\end{bmatrix}+B_{1})...+B_{n-1}\right)+B_{n}$ (26)
@@ 5.3 Thermal inversion
In this example, we aim to model the thermal inversion effect. The numerical results presented in this section are the PINN version of the thermal inversion simulation performed using isogeometric finite element method code [8] described in [12]. The scalar field $u$ in our simulation represents the water vapor forming a cloud. The source represents the evaporation of the cloud evaporation of water particles near the ground. The thermal inversion effect is obtained by introducing the advection field as the gradient of the temperature. Following [13] we define $b(x,y,t)=\frac{\partial T}{\partial y}=-2$ for lower half of the domain ($y<0.5$), and $b(x,y,t)=\frac{\partial T}{\partial y}=2$ for upper half of the domain ($y>0.5$). This represents the thermal inversion effect. The low temperature air mass is kept close to the ground by the higher temperature air mass above it. The temperature gradient decreases from the ground level to half of the domain, representing the low temperature air mass, and then the temperature gradient increases, which represents the temperature increase in the higher temperature air mass. The horizontal diffusion coefficient $K_{x}=0.1$ is stronger than the vertical diffusion coefficient $K_{y}=0.01$. The source function $f(x,y,t)$ represents the cloud vapor that evaporates from the ground level.
@@
$\frac{\partial u(x,y,t)}{\partial t}+\left(b(x,y,t)\cdot\nabla\right)u(x,y,t)-\nabla\cdot\left(K\nabla u(x,y,t)\right)=f(x,y,t),\textrm{ }(x,y,t)\in\Omega\times(0,T]$ (30)
$\frac{\partial u(x,y,t)}{\partial n}=0,\textrm{ }(x,y,t)\in\partial\Omega\times(0,T]$ (31)
$u(x,y,0)=u_{0}(x,y),\textrm{ }(x,y,t)\in\Omega\times 0$ (32)
@@
$\frac{\partial u(x,y,t)}{\partial t}+\frac{\partial T(y)}{\partial y}\frac{\partial u(x,y,t)}{\partial y}-0.1\frac{\partial u(x,y,t)}{\partial x^{2}}-0.01\frac{\partial u(x,y,t)}{\partial y^{2}}=f(x,y,t),\textrm{}(x,y,t)\in\Omega\times(0,T]$ (34)
$\frac{\partial u(x,y,t)}{\partial n}=0,\textrm{ }(x,y,t)\in\partial\Omega\times(0,T]$ (35)
$u(x,y,0)=u_{0}(x,y).\textrm{ }(x,y,t)\in\Omega\times 0$ (36)
@@ 5.4 Tumor growth
$\frac{\partial u(x,y,t)}{\partial n}=0,\textrm{ }(x,y,t)\in\partial\Omega\times(0,T]$ (41)
$u(x,y,0)=u_{0}(x,y),(x,y,t)\in\Omega\times 0$ (42)
@@
$-\frac{\partial D(x,y)}{\partial y}\frac{\partial u(x,y,t)}{\partial y}-D(x,y)\frac{\partial^{2}u(x,y,t)}{\partial y^{2}}-\rho u(x,y,t)\left(1-u(x,y,t)\right)=0$ (44)
Here, $D(x,y)$ represents the tissue density coefficient, where $D(x,y)=0.13$ for the white matter, $D(x,y)=0.013$ for the gray matter, and $D(x,y)=0$ for the cerebrospinal fluid (see [14] for more details). Additionally, $\rho=0.025$ denotes the proliferation rate of the tumor cells. We simplify the model, by removing the derivatives of the tissue density coefficient (which is zero inside a given tissue):
@@
$\left(\frac{\partial u(x,y,t)}{\partial t}-D(x,y)\frac{\partial^{2}u(x,y,t)}{\partial x^{2}}-D(x,y)\frac{\partial^{2}u(x,y,t)}{\partial x^{y}}-\rho u(x,y,t)\left(1-u(x,y,t)\right)\right)^{2}$ (47)